### v1

@@ 6 The memory capacity results
Because the only two properties of self-attention that were used in the proof of Theorem 5 were that (1) $f(z,u,\cdot)$ maps injectively into \$\0\mathbb{R}\backslash\{0\}$ for generic $(z,u)$ and (2) we can reduce to the $d=1$ case, it is natural to wonder what alternatives to self-attention share this property. As it turns out, it works to simply take the weighted average of embedded vectors, a mechanism that we call “token averaging.”
Theorem 6.
Let $\omega,\tau\in\mathbb{N}$. Define
$f:\leavevmode\nobreak$ $\mathbb{R}^{\omega}\times\mathbb{R}^{\tau}\times\cup_{t=1}^{\tau}[\omega]^{t}\to\mathbb{R}$
$(z,u,\alpha)\mapsto\sum_{t=1}^{\mathrm{len}(\alpha)}u_{t}z_{\alpha_{t}}.$
Proof.
First, let $A$ denote $\cup_{t=1}^{\tau}[\omega]^{t}$ and define
Note that the set of $(z,u)$ for which $f(z,u,\cdot)$ does not map injectively into \$\0\mathbb{R}\backslash\{0\}$ is precisely the zero set of $g$. Thus, the result follows if and only if $g$ is nontrivial if and only if (1) each $f(\cdot,\cdot,\alpha)$ is nontrivial and (2) each $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)$ is nontrivial. Note that $f(\mathbbm{1}_{\omega},\mathbbm{1}_{\tau},\alpha)=\mathrm{len}(\alpha)$. This proves (1). It also proves (2) in the case where $\mathrm{len}(\alpha)\neq\mathrm{len}(\beta)$. All that is left is to prove (2) in the case where $\mathrm{len}(\alpha)=\mathrm{len}(\beta)$. Let $\alpha,\beta\in A$ such that $\alpha\neq\beta$ and $\sigma\coloneqq\mathrm{len}(\alpha)=\mathrm{len}(\beta)$. Then there exists $s\in[\sigma]$ such that $\alpha_{s}\neq\beta_{s}$. Choose such an $s$. Set $z=(1,\ldots,\omega)$ and $u=e_{s}$. Then $f(z,u,\alpha)=\alpha_{s}\neq\beta_{s}=f(z,u,\beta)$, proving (2), and so proving the theorem. ∎
Theorem 7.
Let $d,m,\omega,T\in\mathbb{N}$. For all $\theta\coloneqq(Z,u,W,b,V)\in\mathbb{R}^{d\times\omega}\times\mathbb{R}^{T}\times\mathbb{R}^{d\times m}\times\mathbb{R}^{m}\times\mathbb{R}^{m\times\omega}$, define
$h_{\theta}:\cup_{t=1}^{T}[\omega]^{t}$ $\to\mathbb{R}^{\omega}$
$\alpha$ $\mapsto V^{\top}\psi\left(W^{\top}\sum_{t=1}^{\mathrm{len}(\alpha)}u_{t}z_{\alpha_{t}}+b\right)$
where $\psi:\mathbb{R}\to\mathbb{R}$ is real analytic at some $\eta\in\mathbb{R}$ and not a polynomial there. Then the memory capacity of $\{\phi\circ h_{\theta}\mid\theta\}$ is greater than or equal to $m$.
Proof.
Set $n=m$. Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{T}[\omega]^{t}$. Invoking Theorem 6, choose $(z,u)$ such that $f(z,u,\cdot)$ maps injectively into \$\0\mathbb{R}\backslash\{0\}$. Then $[f(z,u,\alpha^{i})]_{i\in[n]}$ has entries which are nonzero and distinct. Set $b=\eta\mathbbm{1}_{m}$. Invoking Theorem 3, choose $w$ such that $\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ has rank $n$. Let $Y\in\mathbb{R}^{\omega\times n}$. Then there exists $V\in\mathbb{R}^{m\times\omega}$ such that $Y=V^{\top}\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$. Choose such a $V$. Define $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d}$ and $W=\mathbbm{1}_{d}w^{\top}/\sqrt{d}$. Then $W^{\top}Z=wz^{\top}$, so $Y=[h_{\theta}(\alpha^{i})]_{i\in[n]}$. Finally, since $\phi(\mathbb{R}^{\omega})=\mathrm{ri}\Delta^{\omega-1}$, the result follows. ∎
@@ 7 Experiment
All experiments are performed on the TinyStories data set (Eldan and Li, 2023), where we take subsets of size 100, 200, 300, 400, 500, 600, 700, 800, 900, 1000 from the data set. These data subsets have 377, 699, 952, 1123, 1459, 1637, 1896, 2117, 2374, 2637 unique contexts, respectively. Note that in our line plot, the minimum threshold varies for each data subset and is computed by taking the entropy lower bound and multiplying it by 0.01 (this helps to account for the fact that larger data sets have larger entropy lower bounds). We train 8 transformer models for each data subset, with $m$ values 4, 8, 16, 32, 64, 128, 256, 512. For simplicity and reduced computation time, we truncate each sequence to have a length of 10. Finally, we use an embedding size of 16 and the sinusoidal positional embedding from Vaswani et al (2017).
@@ 8 Conclusion
We proved upper and lower bounds on the memory capacity of one-layer decoder-only transformers in the setting of next-token prediction. Assuming that the activation function is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable (e.g. arctan, tanh, GELU), we showed that $\Omega(n\omega)$ parameters are necessary to memorize a data set with $n$ unique contexts and a vocabulary of size $\omega$. Assuming that the activation function is real analytic at a point and not a polynomial there (e.g. arctan, tanh, GELU), we showed that $\Omega(n\omega)$ parameters is also sufficient. Furthermore, we provided numerical evidence that the transformer can already be trained to the entropy lower bound in the $\Theta(n\omega)$ parameter regime. Theoretically investigating optimization and generalization in the $\Theta(n\omega)$ parameter regime is left as a future direction.

### v2

@@ 7 Experiment
All experiments are performed on the TinyStories data set (Eldan and Li, 2023), where we take subsets of size 100, 200, 300, 400, 500, 600, 700, 800, 900, 1000 from the data set. These data subsets have 377, 699, 952, 1223, 1459, 1637, 1896, 2117, 2374, 2637 unique contexts, respectively. Note that in our line plot, the minimum threshold varies for each data subset and is computed by taking the entropy lower bound and multiplying it by 0.01 (this helps to account for the fact that larger data sets have larger entropy lower bounds). We train 8 transformer models for each data subset, with $m$ values 4, 8, 16, 32, 64, 128, 256, 512. For simplicity and reduced computation time, we truncate each sequence to have a length of 10. Finally, we use an embedding size of 16 and the sinusoidal positional embedding from Vaswani et al (2017).
@@
8 Token-averaging
In the proof of Theorem 6.5, all that was required of the embedding and self-attention sub-layers was that the composition of them maps injectively into \$\0\mathbb{R}\backslash\{0\}$. This suggests a simple alternative to self-attention. Let $Z\in\mathbb{R}^{d\times\omega}$, $u\in\mathbb{R}^{\mathbb{N}}$, $W\in\mathbb{R}^{d\times m}$, $b\times\mathbb{R}^{m}$, and $V\in\mathbb{R}^{m\times\omega}$. Let $\theta$ denote $(Z,u,W,b,V)$. Let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at a point and not a polynomial there. Define
$h_{\theta}:\cup_{t=1}^{\infty}[\omega]^{t}$ $\to\mathrm{ri}\Delta^{\omega-1}$
$\alpha$ $\mapsto\phi\left(V^{\top}\psi\left(W^{\top}Z[:,\alpha]u[1:|\alpha|]+b\right)\right).$
We call this a token-averaged FNN. Essentially, we have extended the FNN to be defined on $\cup_{t=1}^{\infty}[\omega]^{t}$ through the mapping $\alpha\mapsto Z[:,\alpha]u[1:|\alpha|]$, which we call token-averaging, and which we replaced self-attention with. It turns out this simple model also has optimal next-token prediction capacity, as we now prove.
Theorem 8.1.
Let $\omega,T\in\mathbb{N}$. Define
$f:\leavevmode\nobreak$ $\mathbb{R}^{\omega}\times\mathbb{R}^{T}\times\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}$
$(z,u,\alpha)\mapsto\sum_{t=1}^{|\alpha|}u_{t}z_{\alpha_{t}}.$
Proof.
First, let $A$ denote $\cup_{t=1}^{T}[\omega]^{t}$ and define
Note that the set of $(z,u)$ for which $f(z,u,\cdot)$ does not map injectively into \$\0\mathbb{R}\backslash\{0\}$ is precisely the zero set of $g$. Thus, the result follows if and only if $g$ is nontrivial if and only if (1) each $f(\cdot,\cdot,\alpha)$ is nontrivial and (2) each $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)$ is nontrivial. Note that $f(\mathbbm{1}_{\omega},\mathbbm{1}_{T},\alpha)=|\alpha|$. This proves (1). It also proves (2) in the case where $|\alpha|\neq|\beta|$. All that is left is to prove (2) in the case where $|\alpha|=|\beta|$. Let $\alpha,\beta\in A$ such that $\alpha\neq\beta$ and $\tau\coloneqq|\alpha|=|\beta|$. Then there exists $s\in[\tau]$ such that $\alpha_{s}\neq\beta_{s}$. Choose such an $s$. Set $z=(1,\ldots,\omega)$ and $u=e_{s}$. Then $f(z,u,\alpha)=\alpha_{s}\neq\beta_{s}=f(z,u,\beta)$, proving (2), and so proving the theorem. ∎
Theorem 8.2.
Let $d,m,\omega\in\mathbb{N}$. For all $\theta\coloneqq(Z,u,W,b,V)\in\mathbb{R}^{d\times\omega}\times\mathbb{R}^{\mathbb{N}}\times\mathbb{R}^{d\times m}\times\mathbb{R}^{m}\times\mathbb{R}^{m\times\omega}$, define
$h_{\theta}:\cup_{t=1}^{\infty}[\omega]^{t}$ $\to\mathbb{R}^{\omega}$
$\alpha$ $\mapsto\phi\left(V^{\top}\psi\left(W^{\top}\sum_{t=1}^{|\alpha|}u_{t}z_{\alpha_{t}}+b\right)\right)$
where $\phi$ is the softmax function and $\psi:\mathbb{R}\to\mathbb{R}$ is real analytic at some $\eta\in\mathbb{R}$ and not a polynomial there. Then the next-token prediction capacity of $\{h_{\theta}\mid\theta\}$ is greater than or equal to $m$.
Proof.
Set $n=m$. Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{\infty}[\omega]^{t}$ be distinct and define $T=\max_{i\in[n]}|\alpha^{i}|$. Invoking Theorem 8.1, choose $(z,u)$ such that $[f(z,u,\alpha^{i})]_{i\in[n]}$ has entries which are nonzero and distinct. Invoking Theorem 6.4, choose $w$ such that $\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$ has rank $n$. Let $Y^{\prime}\in(\mathrm{ri}\Delta^{\omega-1})^{n}$. Then there exists $Y\in\mathbb{R}^{\omega\times n}$ such that $\phi(Y)=Y^{\prime}$ where $\phi$ is applied column-wise. Furthermore, there exists $V\in\mathbb{R}^{m\times\omega}$ such that $Y=V^{\top}\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$. So, by defining $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d}$ and $W=\mathbbm{1}_{d}w^{\top}/\sqrt{d}$, we get that $Y=[h_{\theta}(\alpha^{i})]_{i\in[n]}$, completing the proof. ∎
Interestingly, the proof is much simpler than for the transformer model. However, it is important to keep in mind that the memory capacity of a model is only one aspect of it. While the memory capacity is interrelated with the optimization and generalization properties of a model, it does not determine them. Thus, it is important to analyze the models that are actually used in practice, which we have done for transformers. On the other hand, when this analysis suggests alternative models, as it has here, it makes sense to try them in practice. So, we leave it as a future research direction to explore the usefulness of token-averaged FNNs in practice.
@@ 9 Conclusion
We precisely defined the concept of next-token prediction capacity, proved general upper bounds, and proved a lower bound for one-layer multi-head decoder-only transformers. Assuming that the activation function is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable (e.g. arctan, tanh, GELU), we showed that $\Omega(n\omega)$ parameters are necessary to memorize a data set with $n$ unique contexts and a vocabulary of size $\omega$. Assuming that the activation function is real analytic at a point and not a polynomial there (e.g. arctan, tanh, GELU), we showed that $\Omega(n\omega)$ parameters is also sufficient. Furthermore, we provided numerical evidence that the transformer can already be trained to the entropy lower bound in the $\Theta(n\omega)$ parameter regime. Theoretically investigating optimization and generalization in the $\Theta(n\omega)$ parameter regime is left as a future direction. It is also left as a future direction to analyze the next-token prediction capacity for multi-layer transformers and for modified architectures, such as Mixture of Softmaxes, which are not back-loaded, i.e. which have a greater proportion of parameters in the embedding and self-attention sub-layers than in the FNN sub-layer. Finally, it seems worthwhile to consider memory capacity in the more constrained setting of stationary processes.